OpenStudy (briana.img):

A sporting goods store sells sets of golf clubs with steel shafts for $359. Sets of golf clubs with graphite shafts cost $559. Last month the store sold 14 sets of golf clubs for $6,826. How many sets of each type of golf club were sold? Set up and solve a system of equations to find the answer.

OpenStudy (briana.img):

What method is this?

OpenStudy (briana.img):

Does the equation go like this? x+y=14 359x+559y=6826

OpenStudy (briana.img):

@dhirensharma

OpenStudy (anonymous):

yup just a sec....

OpenStudy (anonymous):

tht's right

OpenStudy (anonymous):

go on solve them wat did u get?

OpenStudy (briana.img):

@dhirensharma what method is it?

OpenStudy (anonymous):

to solve system of equation where there are two variables you can either use substitution or elimination

OpenStudy (briana.img):

@dhirensharma oh ok i'll get to solving it

OpenStudy (anonymous):

yup go on

OpenStudy (briana.img):

does -4467y=6467 go well?!

OpenStudy (briana.img):

because i'm getting a fraction when i divide

OpenStudy (anonymous):

no thats not right..

OpenStudy (briana.img):

359x+559y=6826 359(14-y)+559y=6826 359-5026y+559y=6826 -5026y+559Y=6826-359 -4467y=6467

OpenStudy (anonymous):

wrong multiplication

OpenStudy (briana.img):

omg again

OpenStudy (anonymous):

359(14-y)+559y=6826 359(14-y)+559y=6826

OpenStudy (anonymous):

u got -5026y+559Y=6826-359

OpenStudy (anonymous):

try doing multiplication by using distributive property 359*14-359y+559y=6826

OpenStudy (briana.img):

so it'd be 4667?

OpenStudy (briana.img):

4667y

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

5026-359y+559y=6826

OpenStudy (anonymous):

5026+200y=6826

OpenStudy (anonymous):

200y=1800 y=9

OpenStudy (briana.img):

@dhirensharma oh okay thanks!!

OpenStudy (anonymous):

@briana.img sorry i gotta rush but stay in touch if u ever need any help and you can message me a picture of your questions and then i can solve them on a seprate piece of paper and i will take a picture of it and send it to you and then you can look at it whenever you feel like you are having a problem......

OpenStudy (briana.img):

@dhirensharma thank you so much!! you helped me a lot with this exam that i thought i could never do!!

OpenStudy (anonymous):

i will try solve all of them today but just in case i don't i will make sure i send the answers tommorow fr sure bt don't frget to send me the questions or it's picture

OpenStudy (anonymous):

and i will make a video explaining all concepts related o this and i will surely send it to you tommorow fr sure and don't worry you are pretty good at this stuff !!!! ;)

OpenStudy (briana.img):

@dhirensharma good luck on whatever you're doing!!

OpenStudy (anonymous):

thanks and don't frget to send me a photo of those questions and hope to see ya around this forum.....byby